OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 216 in3

OpenStudy (misty1212):

hi!

OpenStudy (misty1212):

it will be the cubed root \[\sqrt[3]{216}=6\]

OpenStudy (displayerror):

Each face of a cube is a square and we know that the sides of a square are all equal. Therefore, all the sides of a cube are also equal. The volume is given by \[Vol = L \times W \times H\] Since all the sides are the same, \[L = W = H = x\] In the above line, I set them equal to a variable 'x' We know the volume is then given by \[Vol = x^3 = 216 \ in^3\] We can then solve for x by using what @misty1212 suggested, which is the cube root.